Mel Bella
Food(F) 60 30
Drinks(D) 30 30

Mel and Bella can produce any linear combination of the production combinations of food or drinks shown in the table. Measuring food on the horizontal axis, their joint production possibility frontier

Select one:

a. will kink away from the origin at 60 units of food.

b. will kink toward the origin at 60 units of food.

c. will kink toward the origin at 30 units of food.

d. will kink away from the origin at 30 units of food.
Answer #1

Opportunity cost of Mel for producing food = 30/60 = 1/2 drinks and of drinks is 2 unitd of food.

Opportunity cost of Bella for producing food is 30/30 = 1 drink and of drinks is 1 unit of food.

Total food that can be produced is 60+30=90

Total drinks = 30+30=60

Mell has a lower opportunity cost in producing food as compared to Bella.

So, the first 60 units of food wiill be produced by Mel.

Now the next 30 units of food will be produced by Bella who has a higher opportinity cost.

The slope for the late half has changed. The slope is equal to the opportunity cost.

Hence the correct option is

a.Will kink away from the origin at 60 units of food.